OpenStudy (tia_sweetheart17):

solve by factoring: 14x=-x^2-45

OpenStudy (anonymous):

get it in standard form first: 14x=-x^2-45 x^2 + 14x + 45 = 0 now factor: x^2 + 14x + 45 = 0 (x + 9)(x + 5) = 0 set each factor = 0: x+9 = 0 x+5 = 0 solve.